OpenStudy (anonymous):

Common brass is a copper and zinc alloy containing 37.0% zinc by mass and having a density of 8.48 g/cm3. A fitting composed of common brass has a total volume of 112.5cm3 . How many atoms of copper and zinc does the fitting contain?

OpenStudy (anonymous):

Source : http://www.edcoogle.com/ and Wiki calculate the mass of the allow: d = m/v ------->8.48 g/ml = x/111.0ml ------> 941.28 grams of the alloy Calculate the mass of both zinc and copper in the sample: .37 x 941.28 = 348.27 grams of zinc .63 x 941.28 = 593.01 grams of copper Calculate the amount of atoms of each using Avogadro's number and the molar mass of zinc and copper from the periodic table, and the known masses in the above equation. 348.27 g Zn x 1 mol Zn/65.409 g x 6.02 x 10 ^23 = 3.2054x 10^24 593.01 g Cu x 1 mol Cu/63.546 x 6.02 x 10 ^23 = 5.6179 x 10^24